Stacked Blocks Speeding Up On A Conveyor Belt-Net
Answer #1

Net Force is given by F = Ma

where a is the acceleration.

therefore the net force on A is : (5 + 3 + 1)Ma = 9Ma = (18M) N

net force on B : (3+3)Ma = (12M) N

C: (3+1)Ma = (8M) N

D: (1 + 3 + 5)Ma = (18M) N

E: (1 + 3)Ma = (8M) N

so in descending order of magnitude of net force,

(18M), (18M), (12M), (8M), (8M)